# A Nonperturbative Toolkit for Quantum Gravity

### v2

@@ 1 Introduction
The Euclidean gravitational path integral has played a central tool in studies of quantum gravity. For example, non-perturbative wormholes in the path integral helped to explain how the entropy of Hawking radiation, ostensibly rising forever, eventually declines to zero consistently with unitarity Penington:2019kki ; Almheiri:2019qdq . Likewise, Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Climent:2024trz ; Balasubramanian:2024yxk constructed an infinite family of black hole microstates under good semiclassical control and showed that their overlaps receive nonpertubative corrections from wormholes. These overlaps led to a Hilbert space of precisely the right dimension to account for the Beckenstein-Hawking entropy.
@@
Strictly speaking, the Euclidean gravitational path integral, understood as a sum over topologies and metrics with a defined boundary condition and weighted by the exponential of the Einstein-Hilbert action, is not well defined by itself. We will understand it as an effective description below some cutoff, completed at high energies by some more refined description, perhaps string theory. However, the gravitational path integral does not seem to define a standard effective field theory. For example, some amplitudes that it computes vanish, even though the square of these amplitudes do not due to non-perturbative wormhole contributions. This suggests an intepretation that the gravitational path integral computes coarse-grained information regarding some underlying fine-grained theory; see, for example, Saad:2019lba ; Saad:2019pqd ; Sasieta:2022ksu ; Marolf:2020xie ; deBoer:2023vsm ; deBoer:2024mqg . Even understood this way, in many settings it is impossible to calculate explicitly, and we resort to approximations like summing over saddlepoints (extrema of the action) of the dominant topologies. Here, we propose a computational toolkit that allows us to extract exact relations in the underlying fine-grained theory despite the coarse-gaining and the difficulty of doing explicit calculations.
@@
The toolkit includes three components building on results in Iliesiu:2024cnh ; Balasubramanian:2024yxk . First, we propose a general way of resolving the identity in the quantum gravity Hilbert space with an overcomplete basis of microstates under semiclassical control. Second, we show that inserting this resolution of the identity into the path integral drastically simplifies the sum over topologies in a limit where the basis is infinitely overcomplete. Third, in this limit, we develop a method for conducting surgery on geometries contributing to the path integral that allows us to demonstrate equality between gravitational path integrals for different quantities, even if neither can be calculated explicitly. As an application we demonstrate that in quantum gravity with two asymptotically locally AdS or flat boundaries the thermal partition function factorizes, even though the sum over geometries includes bulk wormholes. This in turn implies that the trace of any function of the Hamiltonian factorizes. Our computations make no assumptions about holographic duality. Along the way we show that states like the ones used in Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Climent:2024trz ; Balasubramanian:2024yxk to account for the entropy of black holes within different microcanonical windows actually span the complete gravity Hilbert space. As a consequence, a state corresponding to a black hole geometry can be written as a superposition of states that do not contain horizons, but are entangled with a compact universe.
@@ 2 The gravitational path integral
where $\mathcal{M}$ is a bulk manifold filling in the boundary. We have used the Einstein-Hilbert action above, but could have instead used any self-consistent effective theory in any dimension, such as two-dimensional Jackiw-Teitelboim gravity. We understand this path integral as an effective description below some cutoff, completed at high energies by some ultraviolet finite theory. We can construct states of the theory by partitioning the boundary $\mathcal{M}_{b}=\mathcal{M}_{1}\cup\mathcal{M}_{2}$ so that $\partial\mathcal{M}_{1}=\partial\mathcal{M}_{2}=\mathcal{X}$ (see Fig. 1). The path integral then produces a functional of the fields and geometry on the cut $\mathcal{X}$, and formally defines states $|\mathcal{M}_{1,2}\rangle\in\mathcal{H}_{\mathcal{X}}$ in a Hilbert space $\mathcal{H}_{\mathcal{X}}$ such that $\langle\mathcal{M}_{2}|\mathcal{M}_{1}\rangle=\zeta[\Sigma_{b}]$. In general, the path integral cannot be performed exactly, and so we resort to saddlepoint approximations to evaluate overlaps of states constructed in this way.
@@
For a given cut $\mathcal{X}$ one option for producing different states is to fix the boundary manifold $\mathcal{M}_{b}$ and to create states |$\ket{i}$ by inserting matter operators $\mathcal{O}_{i}$ on $\mathcal{M}_{b}$ (see, e.g., Marolf:2017kvq ; Marolf:2020xie ; Colafranceschi:2023moh ). The overlap of such states, $\braket{i}{j}$ is computed by evaluating the path integral with $\mathcal{O}_{i,j}$ inserted on $\mathcal{M}_{1,2}$. Likewise we can calculate the square of the overlap $\braket{i}{j}\braket{j}{i}$ by inserting $\mathcal{O}_{i,j}$ on two copies of the boundary manifold and summing over all Euclidean bulk geometries and topologies that fill in these boundary conditions. Recent work has shown that path integral for square of the overlap can have additional contributions beyond a product of the contributions to $\braket{i}{j}$ – there can be connected topologies, i.e., wormholes, that interpolate between the two boundaries. These non-perturbative contributions, which are akin to instantons in quantum field theory, can produce an apparently paradoxical result where
$\braket{i}{j}=\delta_{ij}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{while}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \braket{i}{j}\braket{j}{i}=\delta^{2}_{ij}+Z_{WH}\,.$ (2)
where $Z_{WH}$ is the wormhole contribution, at least in the semiclassical saddlepoint approximation of the gravitational path integral. Following Saad:2019lba ; Saad:2019pqd ; Sasieta:2022ksu ; Marolf:2020xie ; deBoer:2023vsm ; deBoer:2024mqg ; Penington:2019kki ; Almheiri:2019qdq ; Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Climent:2024trz ; Balasubramanian:2024yxk ; Marolf:2024jze we interpret this situation as arising because the gravitational path integral, treated as an effective theory below some cutoff, and certainly after taking a saddlepoint approximation, calculates coarse-grained observables that compute some kind of average over corresponding quantities in an underlying more fundamental theory that we refer to as the fine-grained theory. To reflect this interpretation we will write an overbar over quantities computed using the gravitational path integral, for example,
@@
Thus, it can transpire that $\overline{\braket{i}{j}}$ vanishes because of cancellation of random phases in the coarse-grained computation of overlaps of states in the fine-grained theory, while these phases cancel in the magnitude squared $\overline{\braket{i}{j}\braket{j}{i}}$. If so, the square root of the latter quantity quantifies the actual magnitude of the overlap in the fine-grained theory.
@@
The Hamiltonian can alternatively viewed as creating a state on two copies of $\mathcal{X}$ – pictorially, we imagine bending the cylinder in Fig. 2(a) around into a half-torus. We denote the state constructed in this manner by the operator $e^{-\frac{\beta H}{2}}$ as $|\beta\rangle$ (note the $1/2$ in the exponent). In QFT, a similar half-torus path integral prepares the Thermofield Double (TFD) state $|TFD(\beta)\rangle=\sum_{i}e^{\frac{-\beta E_{n}}{2}}|n\rangle^{*}_{L}|n\rangle_{R}$ defined on two copies of the theory. In what follows we refer to the norm
$\braket{\beta}{\beta}\equiv Z(\beta)$ (4)
as the “Z partition function”.
@@
Consider a set of $\kappa$ states $\{|i\rangle\}$ constructed as described above by cutting open the path integral with inserted operators. The span of these states, $\mathcal{H}_{\kappa}\equiv span\{|i\rangle,i=i,2,...\kappa\}$ is a subspace of the full Hilbert space. As we discussed, even if these states are orthogonal in the low-energy effective theory, they can have overlaps in the fine-grained theory, arising from nonperturbative effects. We characterize these overlaps by the Gram matrix $G_{ij}\equiv\braket{i}{j}$. Because of these overlaps the dimension of $\mathcal{H}_{\kappa}$ can be less than $\kappa$, the number of states in $\{|i\rangle\}$, i.e., the latter can be overcomplete as a basis for $\mathcal{H}_{\kappa}$. The Gram matrix can be used to construct an orthonormal basis $\{|v_{\gamma}\rangle\}$ from an overcomplete one:
@@
where $U$ is a unitary transformation that diagonalizes $G$ ($G=UDU^{\dagger}$ with $D$ diagonal) (see Appendix A and Appendix A of Balasubramanian:2024yxk for details). Here and in what follows we will use the convention that repeated indices are summed. Following Boruch:2024kvv ; Balasubramanian:2024yxk the inverse of the Gram matrix is defined via analytic continuation of $G^{n}_{ij}$ from positive integers $n$ such that
@@
Thus, even if {|$\{\ket{i}\}$ is overcomplete as a basis for the span, we can use these states and the associated Gram matrix to compute traces of operators. Another useful quantity is the resolution of the identity on $\mathcal{H}_{\kappa}$
@@
$\dim(\mathcal{H}_{\kappa})=Tr_{\mathcal{H}_{\kappa}}(\mathds{1})=G^{-1}_{ij}\braket{j}{i}=\lim_{n\to-1}G^{n}_{ij}\braket{j}{i}=\lim_{n\to-1}\mathrm{Tr}(G^{n+1})$ (10)
@@
Consider quantum gravity with an asymptotic boundary having two disconnected components. We can build states for this theory by cutting open the Euclidean path integral on a slice with two disconnected boundary components $\mathcal{X}=\mathcal{B}_{L}\cup\mathcal{B}_{R}$. We will call this Hilbert space $\mathcal{H}_{L\cup R}$. There are two ways of preparing states in $\mathcal{H}_{L\cup R}$. First, we could fill in and cut open a single connected boundary manifold, as in Fig. 3(a). For example, the state |$\ket{\beta}$ is of this kind. Let us call the resulting Hilbert $\mathcal{H}_{LR}$. Alternatively we could cut open two disconnected closed boundary manifolds $\mathcal{M}_{L}$ and $\mathcal{M}_{R}$ with cuts $\mathcal{B}_{L}$ and $\mathcal{B}_{R}$. As the two boundary manifolds are disconnected, this construction prepares states in a tensor product Hilbert space which we call $\mathcal{H}_{\mathcal{B}_{L}}\otimes\mathcal{H}_{\mathcal{B}_{R}}$ (Fig. 3(b)).
We might expect that the full Hilbert space would be a union $\mathcal{H}_{LR}\cup\mathcal{H}_{\mathcal{B}_{L}}\otimes\mathcal{H}_{\mathcal{B}_{R}}$. But holography with AdS boundary conditions requires that the full Hilbert space with disconnected boundaries has a tensor product structure because the dual CFT Hilbert space does. In other words, we expect that, despite appearances,
@@
In this paper we will take this equality for granted and study the structure of $\mathcal{H}_{LR}$. We will return in a later work to the status of the equality (12).
@@ 3.1 Construction of the shell states
Our candidate basis for the two-sided Hilbert space consists of the shell states studied in Sasieta:2022ksu ; Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Antonini:2023hdh ; see Balasubramanian:2024rek for a comprehensive discussion. We define these states by cutting open the Euclidean gravitational path integral. The resulting boundary $\mathbb{I}\times\mathbb{S}^{d-1}$ has a cut with two boundary components $\mathcal{B}_{L}=\mathcal{B}_{R}$, each of topology $\mathbb{S}^{d-1}$ (Fig. 4). We also insert a $\mathbb{S}^{d-1}$ symmetric heavy ($m\sim\mathit{O}(1/G_{N})$) dust shell operator $\mathcal{O}_{S}$ separated by a boundary length $\frac{\beta_{L}}{2}$ and $\frac{\beta_{R}}{2}$ from the $\mathcal{B}_{L,R}$ cuts. Following Balasubramanian:2022gmo we can obtain an infinite family of shell states in this way by varying the mass $m_{i}$ of the shell operators. We call the resulting state $|i\rangle$.
@@
$|\mathcal{O}\rangle=\sum_{n,m}e^{-(\beta_{L}E_{n}+\beta_{R}E_{m})}\mathcal{O}_{mn}|E_{n},E_{m}\rangle$ (13)
and thus corresponds to a partially entangled state (PETS) Goel:2018ubv . The CFT shell states are defined by a choice of operator $\mathcal{O}$ that is holographically dual to the shell operator insertion on the asymptotic boundary of AdSd+1. In the CFT this is achieved by inserting products of a scalar operator $\mathcal{O}_{\Delta}$ of $\mathit{O}(1)$ scaling dimension $\Delta$. In order to make the states sufficiently heavy $m\sim\mathit{O}(1/G_{N})$, we must insert $n\sim\ell^{d-1}/\mathrm{G_{N}}$ such operators ($\ell$ is the AdS radius) at spherically symmetric points on the $\mathbb{S}^{d-1}$, resulting in the operator $\mathcal{O}_{S}$.
@@
We will want to compute the Gram matrix $G_{ij}\equiv\langle i|j\rangle$ of overlaps between the shell states. We will say that the path integrals calculating these quantities have shell boundaries which include operator insertions $\mathcal{O}_{j}$ and $\mathcal{O}^{\dagger}_{i}$ inserted on the asymptotic boundary separated by asymptotic times $\beta_{L}$ and $\beta_{R}$ respectively (Fig. 5). The overlap $\overline{\langle i|j\rangle}$ is evaluated by the gravitational path integral over geometries satisfying this boundary condition (Fig. 6). Following Balasubramanian:2022gmo , shell states of different mass can be rendered orthogonal in terms of the overlaps computed in Fig. 6 by taking their inertial mass differences to be arbitrarily large, as it takes $|m_{i}-m_{j}|$ bulk interactions in Planck units to match such shells in the bulk, leading to
@@
Figure 7: Construction of shell wormhole saddles $Z_{2}$. We glue two disks together by discarding the regions indicated in purple and attaching the remainder along the shell worldvolumes (red, black dashes indicate the glued shells) while obeying the Israel junction conditions. Saddlepoints, i.e., solutions to the equation of motion, exist with this topology Balasubramanian:2022gmo .

### v3

@@ 1 Introduction
The Euclidean gravitational path integral has played a central tool in studies of quantum gravity. For example, non-perturbative wormholes in the path integral helped to explain how the entropy of Hawking radiation, ostensibly rising forever, eventually declines to zero consistently with unitarity Penington et al. (2022); Almheiri et al. (2020). Likewise, Balasubramanian et al. (2024c, b); Climent et al. (2024); Balasubramanian et al. (2025) constructed an infinite family of black hole microstates under good semiclassical control and showed that their overlaps receive nonpertubative corrections from wormholes. These overlaps led to a Hilbert space of precisely the right dimension to account for the Beckenstein-Hawking entropy.
@@
Strictly speaking, the Euclidean gravitational path integral, understood as a sum over topologies and metrics with a defined boundary condition and weighted by the exponential of the Einstein-Hilbert action, is not well defined by itself. We will understand it as an effective description below some cutoff, completed at high energies by some more refined description, perhaps string theory. However, the gravitational path integral does not seem to define a standard effective field theory. For example, some amplitudes that it computes vanish, even though the square of these amplitudes do not due to non-perturbative wormhole contributions. This suggests an intepretation that the gravitational path integral computes coarse-grained information regarding some underlying fine-grained theory; see, for example, Saad et al. (2019); Saad (2019); Sasieta (2023); Marolf and Maxfield (2020); de Boer et al. (2024a, b). Even understood this way, in many settings it is impossible to calculate explicitly, and we resort to approximations like summing over saddlepoints (extrema of the action) of the dominant topologies. Here, we propose a computational toolkit that allows us to extract exact relations in the underlying fine-grained theory despite the coarse-gaining and the difficulty of doing explicit calculations.
@@
The toolkit includes three components building on results in Iliesiu et al. (2024); Balasubramanian et al. (2025). First, we propose a general way of resolving the identity in the quantum gravity Hilbert space with an overcomplete basis of microstates under semiclassical control. Second, we show that inserting this resolution of the identity into the path integral drastically simplifies the sum over topologies in a limit where the basis is infinitely overcomplete. Third, in this limit, we develop a method for conducting surgery on geometries contributing to the path integral that allows us to demonstrate equality between gravitational path integrals for different quantities, even if neither can be calculated explicitly. As an application we demonstrate that in quantum gravity with two asymptotically locally AdS or flat boundaries the thermal partition function factorizes, even though the sum over geometries includes bulk wormholes. This in turn implies that the trace of any function of the Hamiltonian factorizes. Our computations make no assumptions about holographic duality. Along the way we show that states like the ones used in Balasubramanian et al. (2024c, b); Climent et al. (2024); Balasubramanian et al. (2025) to account for the entropy of black holes within different microcanonical windows actually span the complete gravity Hilbert space. As a consequence, a state corresponding to a black hole geometry can be written as a superposition of states that do not contain horizons, but are entangled with a compact universe.
@@ 2 The gravitational path integral
where $\mathcal{M}$ is a bulk manifold filling in the boundary. We have used the Einstein-Hilbert action above, but could have instead used any self-consistent effective theory in any dimension, such as two-dimensional Jackiw-Teitelboim (JT) gravity. We understand this path integral as an effective description below some cutoff, completed at high energies by some ultraviolet finite theory. We can construct states of the theory by partitioning the boundary $\mathcal{M}_{b}=\mathcal{M}_{1}\cup\mathcal{M}_{2}$ so that $\partial\mathcal{M}_{1}=\partial\mathcal{M}_{2}=\mathcal{X}$ (see Fig. 1). The path integral then produces a functional of the fields and geometry on the cut $\mathcal{X}$, and formally defines states $|\mathcal{M}_{1,2}\rangle\in\mathcal{H}_{\mathcal{X}}$ in a Hilbert space $\mathcal{H}_{\mathcal{X}}$ such that $\langle\mathcal{M}_{2}|\mathcal{M}_{1}\rangle=\zeta[\Sigma_{b}]$. In general, the path integral cannot be performed exactly, and so we resort to saddlepoint approximations to evaluate overlaps of states constructed in this way. Note that defining a state by partitioning boundary data is a standard way of producing states in the bulk quantum gravitational Hilbert space. For instance, such methods were used in Iliesiu et al. (2024) and Marolf and Maxfield (2020) to construct states in JT gravity and a topological toy model respectively, and in this work we will apply this method in higher dimensional Einstein-Hilbert gravity. Crucially, we do not assume any kind of dual boundary CFT construction. It is clear our states live in the bulk Hilbert space as we can, in principle, compute the wavefunctionals $\Psi_{\mathcal{M}_{1}}(h,\phi)\equiv\braket{h,\phi|\mathcal{M}_{1}}$ between our states and states defined by some bulk time-slice data $h_{\mu\nu},\phi$ by performing the path integral subject to these joint boundary conditions.
@@
$\braket{i|j}=\delta_{ij}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathrm{while}\penalty 10000\ \penalty 10000\ \penalty 10000\ \braket{i|j}\braket{j|i}=\delta^{2}_{ij}+Z_{WH}\,.$ (2)
where $Z_{WH}$ is the wormhole contribution, at least in the semiclassical saddlepoint approximation of the gravitational path integral. Following Saad et al. (2019); Saad (2019); Sasieta (2023); Marolf and Maxfield (2020); de Boer et al. (2024a, b); Penington et al. (2022); Almheiri et al. (2020); Balasubramanian et al. (2024c, b); Climent et al. (2024); Balasubramanian et al. (2025); Marolf (2025) we interpret this situation as arising because the gravitational path integral, treated as an effective theory below some cutoff, and certainly after taking a saddlepoint approximation, calculates coarse-grained observables that compute some kind of average over corresponding quantities in an underlying more fundamental theory that we refer to as the fine-grained theory. To reflect this interpretation we will write an overbar over quantities computed using the gravitational path integral, for example,
$\overline{\braket{i_{1}|j_{i}}\braket{i_{2}|j_{2}}\cdots\braket{i_{n}|j_{n}}}\,.$ (3)
Thus, it can transpire that $\overline{\braket{i|j}}$ vanishes because of cancellation of random phases in the coarse-grained computation of overlaps of states in the fine-grained theory, while these phases cancel in the magnitude squared $\overline{\braket{i|j}\braket{j|i}}$. If so, the square root of the latter quantity quantifies the actual magnitude of the overlap in the fine-grained theory.
@@
The Hamiltonian can alternatively viewed as creating a state on two copies of $\mathcal{X}$ – pictorially, we imagine bending the cylinder in Fig. 2(a) around into a half-torus. We denote the state constructed in this manner by the operator $e^{-\frac{\beta H}{2}}$ as $|\beta\rangle$ (note the $1/2$ in the exponent). In QFT, a similar half-torus path integral prepares the Thermofield Double (TFD) state $|TFD(\beta)\rangle=\sum_{n}e^{\frac{-\beta E_{n}}{2}}|E_{n}\rangle^{*}_{L}|E_{n}\rangle_{R}$ defined on two copies of the theory. In what follows we refer to the norm
$\braket{\beta|\beta}\equiv Z(\beta)$ (4)
as the “Z partition function”. We introduce this terminology as it is not clear from the gravitational Hilbert space perspective presented here why this path integral should compute the thermal partition sum of the theory. In Balasubramanian and Yildirim (2025a) we use the toolkit developed here to show that this is indeed the case, once non-perturbative wormhole corrections to the thermal trace are taken into account.
@@
Consider a set of $\kappa$ states $\{|i\rangle\}$ constructed as described above by cutting open the path integral with inserted operators. The span of these states, $\mathcal{H}_{\kappa}\equiv span\{|i\rangle,i=i,2,...\kappa\}$ is a subspace of the full Hilbert space. As we discussed, even if these states are orthogonal in the low-energy effective theory, they can have overlaps in the fine-grained theory, arising from nonperturbative effects. We characterize these overlaps by the Gram matrix $G_{ij}\equiv\braket{i|j}$. Because of these overlaps the dimension of $\mathcal{H}_{\kappa}$ can be less than $\kappa$, the number of states in $\{|i\rangle\}$, i.e., the latter can be overcomplete as a basis for $\mathcal{H}_{\kappa}$. The Gram matrix can be used to construct an orthonormal basis $\{|v_{\gamma}\rangle\}$ from an overcomplete one:
@@
where $U$ is a unitary transformation that diagonalizes $G$ ($G=UDU^{\dagger}$ with $D$ diagonal) (see Appendix A and Appendix A of Balasubramanian et al. (2025) for details). Here and in what follows we will use the convention that repeated indices are summed. Following Boruch et al. (2024a); Balasubramanian et al. (2025) the inverse of the Gram matrix is defined via analytic continuation of $G^{n}_{ij}$ from positive integers $n$ such that
@@
[…]
@@
$\dim(\mathcal{H}_{\kappa})=Tr_{\mathcal{H}_{\kappa}}(\mathds{1})=G^{-1}_{ij}\braket{j|i}=\lim_{n\to-1}G^{n}_{ij}\braket{j|i}=\lim_{n\to-1}\mathrm{Tr}(G^{n+1})$ (10)
@@
[…]
We might expect that the full Hilbert space would be an (internal) direct sum $\mathcal{H}_{LR}\oplus\mathcal{H}_{\mathcal{B}_{L}}\otimes\mathcal{H}_{\mathcal{B}_{R}}$ or perhaps $\mathcal{H}_{L\cup R}=\{u+v\;|\;u\in\mathcal{H}_{LR},\,v\in\mathcal{H}_{\mathcal{B}_{L}}\otimes\mathcal{H}_{\mathcal{B}_{R}}\}$. But holography with AdS boundary conditions requires that the full Hilbert space with disconnected boundaries has a tensor product structure because the dual CFT Hilbert space does. In other words, we expect that, despite appearances,
@@
In this paper we will take this equality for granted and study the structure of $\mathcal{H}_{LR}$. In Balasubramanian and Yildirim (2025b) we show that the factorisation (12) results from the existence of novel Euclidean wormhole saddles connecting these various boundary conditions.
@@ 3.1 Construction of the shell states
Our candidate basis for the two-sided Hilbert space consists of the shell states studied in Sasieta (2023); Balasubramanian et al. (2024c, b); Antonini et al. (2023); see Balasubramanian et al. (2024a) for a comprehensive discussion. We define these states by cutting open the Euclidean gravitational path integral. The resulting boundary $\mathbb{I}\times\mathbb{S}^{d-1}$ has a cut with two boundary components $\mathcal{B}_{L}=\mathcal{B}_{R}$, each of topology $\mathbb{S}^{d-1}$ (Fig. 4). We also insert a $\mathbb{S}^{d-1}$ symmetric heavy ($m\sim\mathit{O}(1/G_{N})$) dust shell operator $\mathcal{O}_{S}$ separated by a boundary length $\frac{\beta_{L}}{2}$ and $\frac{\beta_{R}}{2}$ from the $\mathcal{B}_{L,R}$ cuts. Following Balasubramanian et al. (2024c) we can obtain an infinite family of shell states in this way by varying the mass $m_{i}$ of the shell operators. We call the resulting state $|i\rangle$.
@@
$|\mathcal{O}\rangle=\sum_{n,m}e^{-(\beta_{L}E_{n}+\beta_{R}E_{m})}\mathcal{O}_{mn}|E_{n}\rangle_{L}|E_{m}\rangle_{R}$ (13)
and thus corresponds to a partially entangled state (PETS) Goel et al. (2019). The CFT shell states are defined by a choice of operator $\mathcal{O}$ that is holographically dual to the shell operator insertion on the asymptotic boundary of AdSd+1. In the CFT this is achieved by inserting products of a scalar operator $\mathcal{O}_{\Delta}$ of $\mathit{O}(1)$ scaling dimension $\Delta$. In order to make the states sufficiently heavy $m\sim\mathit{O}(1/G_{N})$, we must insert $n\sim\ell^{d-1}/\mathrm{G_{N}}$ such operators ($\ell$ is the AdS radius) at spherically symmetric points on the $\mathbb{S}^{d-1}$, resulting in the operator $\mathcal{O}_{S}$.
To be clear, a particular microstate of the gravity theory is constructed by specifying the microscopic details of the operator insertions that make up the heavy shell. There are many such exact microstates which share the same semiclassical description as a spherically symmetric heavy dust shell in the bulk. As discussed Sec. 2 above, the semiclassical gravitational path integral in the manner used here appears to compute a coarse-graining over these fine-grained details. Thus while we have some particular microscopic realization in mind for each of our shell states, we can, remarkably, learn much about the underlying theory without specifying these details.
@@
We will want to compute the Gram matrix $G_{ij}\equiv\langle i|j\rangle$ of overlaps between the shell states. We will say that the path integrals calculating these quantities have shell boundaries which include operator insertions $\mathcal{O}_{j}$ and $\mathcal{O}^{\dagger}_{i}$ inserted on the asymptotic boundary separated by asymptotic times $\beta_{L}$ and $\beta_{R}$ respectively (Fig. 5). The overlap $\overline{\langle i|j\rangle}$ is evaluated by the gravitational path integral over geometries satisfying this boundary condition (Fig. 6). Following Balasubramanian et al. (2024c), shell states of different mass can be rendered orthogonal in terms of the overlaps computed in Fig. 6 by taking their inertial mass differences to be arbitrarily large, as it takes $|m_{i}-m_{j}|$ bulk interactions in Planck units to match such shells in the bulk, leading to
@@
Figure 7: Construction of shell wormhole saddles $Z_{2}$. We glue two disks together by discarding the regions indicated in purple and attaching the remainder along the shell worldvolumes (red, black dashes indicate the glued shells) while obeying the Israel junction conditions. Saddlepoints, i.e., solutions to the equation of motion, exist with this topology Balasubramanian et al. (2024c).